(1) y+2y+y=x2-1-3, y(0)=-2, y(0)=1 (2) + y= -8cos2x+6sin 2 x (3) y*- 3y + 2 y =e (x2 + 2x - 1)

D.E
Answer #1

3) 닝 - 34 +3j e fx+ 2 -1) y -3Dy 424= (x+2) 0- 30t2) y e (x+2) e m를 3m4220 m2 2m -mt220. m(m-29-1 (R-20 (m-9{m-2) =os m1,2 C